# Quantifying Interpretability in CLIP Models with Concept Consistency

### v1

@@ Abstract
CLIP is one of the most popular foundational models and is heavily used for many vision-language tasks. However, little is known about the inner workings of CLIP. While recent work has proposed decomposition-based interpretability methods for identifying textual descriptions of attention heads in CLIP, the implications of conceptual consistency in these text labels on interpretability and model performance has not been explored. To bridge this gap, we study the conceptual consistency of text descriptions for attention heads in CLIP-like models. We conduct extensive experiments on six different models from OpenAI and OpenCLIP which vary by size, type of pre-training data and patch size. We propose Concept Consistency Score (CCS), a novel interpretability metric that measures how consistently individual attention heads in CLIP models align with specific concepts. To assign concept labels to heads, we use in-context learning with ChatGPT, guided by a few manually-curated examples, and validate these labels using an LLM-as-a-judge approach. Our soft-pruning experiments reveal that high CCS heads are critical for preserving model performance, as pruning them leads to a significantly larger performance drop than pruning random or low CCS heads. Notably, we find that high CCS heads capture essential concepts and play a key role in out-of-domain detection, concept-specific reasoning, and video-language understanding. These results position CCS as a powerful interpretability metric for analyzing CLIP-like models.
@@ 1 Introduction
Large-scale vision-language (VL) models such as CLIP  [30] have significantly advanced state-of-the-art performance in vision tasks in recent years. Consequently, CLIP has been extensively used as a foundational model for downstream tasks such as video retrieval, image generation, and segmentation  [26, 25, 4, 13, 22]. This has enabled the construction of compositional models combining CLIP with other foundation models, thereby increasing the functionality of CLIP while also adding complexity to the overall model structure. Furthermore, tools and methods for interpreting the functioning of such models are becoming increasingly important in order to enable their adoption in domains where an understanding of model predictions is required. Therefore, given its widespread use as a foundation model for VL tasks, it is imperative to understand the inner workings of CLIP models.
Recently, Gandelsman et al. [15] proposed an interpretability approach for CLIP-like models which decomposes image representations into a sum over contributions from individual attention heads, layers, and image patches. This enables the identification of text sequences which characterize the semantic role played by different attention heads of CLIP. While this provides valuable insights into the inner workings of CLIP, a critical question remains regarding the conceptual consistency of text labels automatically identified for various attention heads of the model. This question has important implications for the interpretability, as lower levels of consistency in the concepts identified for any individual attention head could make its functional role in the model more difficult to understand.
@@
To address this important question, we introduce a methodology for systematically measuring the conceptual consistency of text descriptions extracted for describing the semantic roles of attention heads in CLIP-like models. We conduct our study using six different models: ViT-B-16, ViT-B-32, and ViT-L-14 from OpenAI, as well as ViT-B-16, ViT-B-32, and ViT-L-14 from OpenCLIP [21]. First, we identify interpretable structures within the individual heads of the last four layers of the model using a set of text descriptions. To accomplish this, we employ the TextSpan algorithm  [15], which helps us find the most appropriate text descriptions for each head. After identifying these text descriptions, we assign labels to each head representing the common property shared by the descriptions. This labeling process is carried out using in-context learning with ChatGPT. We begin by manually labeling five pairs of text descriptions and their corresponding property labels, which serve as examples. These examples are then used to prompt ChatGPT to assign labels for the remaining heads.
@@
Leveraging the resulting text descriptions of attention heads, we introduce the Concept Consistency Score (CCS), a new interpretability metric that quantifies how strongly individual attention heads in CLIP models align with specific concepts. Using GPT-4o as an automatic judge, we compute CCS for each head and classify them into high, moderate, and low categories based on defined thresholds. A key contribution of our work is our targeted soft-pruning experiments which show that heads with high CCS are essential for maintaining model performance; pruning these heads causes a significantly larger performance drop compared to pruning random or low CCS heads. We also show that high CCS heads are not only crucial for general vision-language tasks but are especially important for out-of-domain detection and concept-specific reasoning. Additionally, our experiments in video retrieval highlight that high CCS heads are equally vital for temporal and cross-modal understanding, thereby underscoring the broad relevance of CCS in analyzing and interpreting CLIP-like models.
@@ 2 Related Work
Early research on interpretability primarily concentrated on convolutional neural networks (CNNs) due to their intricate and opaque decision-making processes [37, 32, 33, 14, 18]. More recently, the interpretability of Vision Transformers (ViT) has garnered significant attention as these models, unlike CNNs, rely on self-attention mechanisms rather than convolutions. Researchers have focused on task-specific analyses in areas such as image classification, captioning, and object detection to understand how ViTs process and interpret visual information [10, 12, 28, 36, 9, 11]. One of the key metrics used to measure interpretability in ViTs is the attention mechanism itself, which provides insights into how the model distributes focus across different parts of an image when making decisions [8, 6]. This has led to the development of techniques that leverage attention maps to explain ViT predictions. Early work on multimodal interpretability, which involves models that handle both visual and textual inputs, probed tasks such as how different modalities influence model performance [5, 27] and how visual semantics are represented within the model [17, 24]. Aflalo et al. [1] explored interpretability methods for vision-language transformers, examining how these models combine visual and textual information to make joint decisions. Similarly, Stan et al. [34] proposed new approaches for interpreting vision-language models, focusing on the interactions between modalities and how these influence model predictions. Our work builds upon and leverages the methods introduced by Gandelsman et al. [15, 16] to interpret attention heads, neurons, and layers in vision-language models, providing deeper insights into their decision-making processes.
Model Model size Patch size Pre-training data Kappa Spearman’s correlation ($\rho$) Kendall ($\uptau$)
CLIP B 32 OpenAI-400M 0.757 0.781 0.781
CLIP B 16 OpenCLIP-LAION2B 0.676 0.678 0.678
CLIP L 14 OpenAI-400M 0.758 0.758 0.758
@@
Table 2: Results between human judgment and LLM judgment on CCS labelling.
Model Model size Patch size Pre-training data CCS@0 CCS@1 CCS@2 CCS@3 CCS@4 CCS@5
CLIP B 32 OpenAI-400M 0.021 0.062 0.167 0.271 0.312 0.167
CLIP B 32 OpenCLIP-datacomp 0.104 0.062 0.208 0.189 0.208 0.229
CLIP B 16 OpenAI-400M 0.021 0.062 0.125 0.292 0.292 0.208
CLIP B 16 OpenCLIP-LAION2B 0.062 0.062 0.105 0.25 0.25 0.271
CLIP L 14 OpenAI-400M 0.062 0.109 0.172 0.204 0.203 0.25
CLIP L 14 OpenCLIP-LAION2B 0.016 0.031 0.109 0.219 0.297 0.328
Table 3: Concept Consistency Score (CCS) for CLIP models.
3 Methodology
In this section, we outline the methodology used in our analysis, beginning with an explanation of the TextSpan algorithm. We then describe how we extend this algorithm to apply it across all attention heads in multiple CLIP models using in-context learning.
The TextSpan algorithm  [15] is designed to decompose individual attention heads by associating them with corresponding text descriptions. It requires an initial set of text descriptions broadly capturing the concepts in an image. The algorithm starts by generating two matrices: $C$, which contains the outputs for a specific head (denoted as $(l,h)$), and $R$, which includes representations of candidate text descriptions projected onto the span of $C$. In each iteration, TextSpan calculates the dot product between each row of $R$ and the outputs in $C$ to identify the row with the highest variance, known as the "first principal component." Once identified, this component is projected away from all rows, and the process is repeated to find additional components. The projection step ensures that each new component adds variance that is orthogonal to the earlier components, thereby isolating distinct aspects of the text descriptions relevant to each head.
TextSpan is effective at identifying text descriptions that are most similar to a given head. Note that, TextSpan associates only the text descriptions to a head and doesn’t label the common concept. In the next section, we explain in detail our approach to label the concept learnt by individual heads in CLIP models.
@@ 4.1.3 Evaluating Concept Consistency
To assess the consistency of a head with respect to its assigned concept label, we employ a state-of-the-art vision-language model, GPT-4o, as an external evaluator. For each TextSpan $T_{i}$ associated with head $h$, GPT-4o determines whether it aligns with the assigned concept $C_{h}$. The Concept Consistency Score (CCS) for head $h$ is then computed as:
@@
where $\mathbb{1}[\cdot]$ is an indicator function that returns 1 if GPT-4o judges $T_{i}$ to be consistent with $C_{h}$, and 0 otherwise.
@@
By computing CCS across all attention heads and layers in a CLIP model, we can systematically interpret how different heads specialize in capturing semantic concepts. This provides crucial insights into the internal representations learned by CLIP, revealing whether specific concepts (e.g., object categories, attributes, locations) are localized within certain heads or distributed across multiple layers. This approach advances model interpretability by offering a quantifiable and automated means to analyze the emergent concepts within CLIP’s attention mechanisms, bridging the gap between learned representations and human-understandable concepts.
@@ 4.2 Evaluating LLM Judgment Alignment with Human Annotations
In the previous section, we introduced the Concept Consistency Score (CCS), which is computed using GPT-4o as an external evaluator. However, this reliance on a large language model (LLM) for judgment raises a critical question: Are LLM evaluations reliable and aligned with human assessments? To address this, we conduct a human evaluation study to systematically compare LLM judgments with human annotations.
To assess the reliability of LLM-based evaluations, we collect 50 TextSpan descriptions from three different models, along with their corresponding concept labels assigned during CCS computation. One of the authors of this work then evaluated whether each TextSpan is semantically related to its assigned concept label. The collected human ratings serve as the ground truth against which LLM judgments are compared. To measure the alignment between human and LLM evaluations, we compute the following statistical agreement metrics: Cohen’s Kappa, Spearman’s Rank Correlation ($\rho$) and Kendall’s Tau ($\uptau$).
Table 2 presents the computed agreement scores. Notably, the Cohen’s Kappa values exceed 0.65, indicating substantial agreement between human and LLM evaluations. Similarly, the Spearman’s correlation and Kendall’s Tau values are consistently above 0.65, demonstrating a positive correlation between LLM and human judgments. These results provide strong empirical support for using LLMs as evaluators in concept consistency analysis. The high alignment between human and LLM judgments indicates that LLMs can reliably assess semantic coherence, reducing the need for costly and time-consuming human annotations. In the next section, we introduce tasks and the corresponding datasets used in our experiments.
4.3 Experimental Setting
4.3.1 Tasks
We primarly focus our evaluation on image classification, out-of-domain classification and video retrieval tasks. Below, we list the evaluation datasets for each of the tasks.
Image classification: CIFAR-10  [23], CIFAR-100  [23], Food-101  [3], Country-211  [30] and Oxford-pets  [29].
Out-of-domain classification: Imagenet-A  [20] and Imagenet-R  [19].
Video retrieval: MSRVTT  [35], MSVD  [7], DiDeMo  [2].
4.3.2 Models
For experiments we use the following six foundational image-text models: ViT-B-32, ViT-B-16 and ViT-L-14 pretrained from OpenAI-400M  [30] and LAION2B  [31]. Next, we discuss in detail the results from the experiments.
@@ 4.3.2 Models
Model Country-211 Oxford-pets
Original High CCS Low CCS Original High CCS Low CCS
ViT-B-32-OpenAI 17.16 11.46 16.3 50.07 46.66 48.96
ViT-B-32-datacomp 4.43 4.37 4.37 26.48 25.98 25.33
ViT-B-16-OpenAI 22.81 10.72 21.79 52.72 49.12 51.89
ViT-B-16-LAION 20.45 7.49 16.87 65.79 48.48 49.81
ViT-L-14-OpenAI 31.91 23.21 30.63 61.79 62.04 62.08
ViT-L-14-LAION 26.41 16.38 25.66 54.1 56.12 57.16
Table 5: Accuracy comparison of various CLIP models on Country-211 and Oxford-pets datasets. The values represent original accuracy, performance after pruning high-CCS heads, and performance after pruning low-CCS heads.
@@
Figure 2: Zero-shot performance comparison for CIFAR-10, CIFAR-100, and Food-101 datasets under different pruning strategies. For random pruning, results are averaged across three runs.
Model Imagenet-A Imagenet-R
Original High CCS Low CCS Original High CCS Low CCS
ViT-B-32-OpenAI 31.49 20.24 28.72 69.09 54.47 64.45
ViT-B-32-datacomp 4.96 4.59 4.65 34.06 31.6 32.47
ViT-B-16-OpenAI 49.85 25.49 47.27 77.37 55.52 74.84
ViT-B-16-LAION 37.97 25.27 27.44 80.56 66.32 71.73
ViT-L-14-OpenAI 70.4 68.15 69.2 87.87 86.56 86.97
ViT-L-14-LAION 53.8 42.44 52.93 87.12 82.22 86.94
Table 6: Accuracy comparison of various CLIP models on ImageNet-A and ImageNet-R. The values represent original accuracy, performance after pruning high-CCS heads, and performance after pruning low-CCS heads.

### v2

# Pruning the Paradox: How CLIP’s Most Informative Heads Enhance Performance While Amplifying Bias

@@ Abstract
CLIP is one of the most popular foundational models and is heavily used for many vision-language tasks. However, little is known about the inner workings of CLIP. While recent work has proposed decomposition-based interpretability methods for identifying textual descriptions of attention heads in CLIP, the implications of conceptual consistency in these text labels on interpretability and model performance has not been explored. To bridge this gap, we study the conceptual consistency of text descriptions for attention heads in CLIP-like models. We conduct extensive experiments on six different models from OpenAI and OpenCLIP which vary by size, type of pre-training data and patch size. We propose Concept Consistency Score (CCS), a novel interpretability metric that measures how consistently individual attention heads in CLIP models align with specific concepts. To assign concept labels to heads, we use in-context learning with ChatGPT, guided by a few manually-curated examples, and validate these labels using an LLM-as-a-judge approach. Our soft-pruning experiments reveal that high CCS heads are critical for preserving model performance, as pruning them leads to a significantly larger performance drop than pruning random or low CCS heads. Notably, we find that high CCS heads capture essential concepts and play a key role in out-of-domain detection, concept-specific reasoning, and video-language understanding. Moreover, we prove that high CCS heads learn spurious correlations amplifying social biases. These results position CCS as a powerful interpretability metric exposing the paradox of performance and social biases in CLIP models.
Machine Learning, ICML
@@ 1 Introduction
Large-scale vision-language (VL) models such as CLIP  (Radford et al., 2021) have significantly advanced state-of-the-art performance in vision tasks in recent years. Consequently, CLIP has been extensively used as a foundational model for downstream tasks such as video retrieval, image generation, and segmentation  (Luo et al., 2022; Liu et al., 2024; Brooks et al., 2023; Esser et al., 2024; Kirillov et al., 2023). This has enabled the construction of compositional models combining CLIP with other foundation models, thereby increasing the functionality of CLIP while also adding complexity to the overall model structure. However, as these models gain prominence in real-world applications, their embedded social biases  (Howard et al., 2024; Hall et al., 2023; Seth et al., 2023) have emerged as a critical concern with potentially harmful downstream consequences. Despite the growing body of work documenting these biases, a fundamental question remains unanswered: what mechanisms within these models’ architectures drive both their impressive capabilities and problematic shortcomings?
Recent interpretability advances  (Gandelsman et al., ) have made initial progress by decomposing CLIP’s image representations into contributions from individual attention heads, identifying text sequences that characterize different heads’ semantic roles. However, this approach provides only a partial view into CLIP’s inner workings, leaving a critical missing piece: systematic understanding of the visual concepts encoded at the attention head level—and how these concepts underpin both the model’s strengths and its social failures.
@@
Our work addresses this critical gap through a novel interpretability framework we call “conceptual consistency”. This framework systematically analyzes which visual concepts are learned by individual attention heads and how consistently these concepts are processed throughout the model’s architecture. First, we identify interpretable structures within the individual heads of the last four layers of the model using a set of text descriptions. To accomplish this, we employ the TextSpan algorithm  (Gandelsman et al., ), which helps us find the most appropriate text descriptions for each head. After identifying these text descriptions, we assign labels to each head representing the common property shared by the descriptions. This labeling process is carried out using in-context learning with ChatGPT. We begin by manually labeling five pairs of text descriptions and their corresponding concept labels, which serve as examples. These examples are then used to prompt ChatGPT to assign labels for the remaining heads.
@@
Leveraging the resulting text descriptions and concept labels of attention heads, we introduce the Concept Consistency Score (CCS), a new interpretability metric that quantifies how strongly individual attention heads in CLIP models align with specific concepts. Using GPT-4o, Gemini and Claude as automatic judges, we compute CCS for each head and classify them into high, moderate, and low categories based on defined thresholds. A key contribution of our work is our targeted soft-pruning experiments which show that heads with high CCS are essential for maintaining model performance; pruning these heads causes a significantly larger performance drop compared to pruning any other heads. We also show that high CCS heads are not only crucial for general vision-language tasks but are especially important for out-of-domain detection and targetted concept-specific reasoning. Additionally, our experiments in video retrieval highlight that high CCS heads are equally vital for temporal and cross-modal understanding. Moreover, we demonstrate that high CCS heads often encode spurious correlations, contributing to social biases in CLIP models. Selective pruning of these heads can reduce such biases without the need for fine-tuning. Together, these results expose a fundamental paradox: while high-CCS heads are indispensable for strong model performance, they are simultaneously key contributors to undesirable biases.
@@ 2 Related Work
Early research on interpretability primarily concentrated on convolutional neural networks (CNNs) due to their intricate and opaque decision-making processes (Zeiler & Fergus, 2014; Selvaraju et al., 2017; Simonyan et al., 2014; Fong & Vedaldi, 2017; Hendricks et al., 2016). More recently, the interpretability of Vision Transformers (ViT) has garnered significant attention as these models, unlike CNNs, rely on self-attention mechanisms rather than convolutions. Researchers have focused on task-specific analyses in areas such as image classification, captioning, and object detection to understand how ViTs process and interpret visual information (Dong et al., 2022; Elguendouze et al., 2023; Mannix & Bondell, 2024; Xue et al., 2022; Cornia et al., 2022; Dravid et al., 2023). One of the key metrics used to measure interpretability in ViTs is the attention mechanism itself, which provides insights into how the model distributes focus across different parts of an image when making decisions (Cordonnier et al., 2019; Chefer et al., 2021). This has led to the development of techniques that leverage attention maps to explain ViT predictions. Early work on multimodal interpretability, which involves models that handle both visual and textual inputs, probed tasks such as how different modalities influence model performance (Cao et al., 2020; Madasu & Lal, 2023) and how visual semantics are represented within the model (Hendricks & Nematzadeh, 2021; Lindström et al., 2021). Aflalo et al. (Aflalo et al., 2022) explored interpretability methods for vision-language transformers, examining how these models combine visual and textual information to make joint decisions. Similarly, Stan et al. (Stan et al., 2024) proposed new approaches for interpreting vision-language models, focusing on the interactions between modalities and how these influence model predictions. Our work builds upon and leverages the methods introduced by Gandelsman et al. ; Gandelsman et al. (2024) to interpret attention heads, neurons, and layers in vision-language models, providing deeper insights into their decision-making processes.
Model Kappa SC ($\rho$) Kendall ($\uptau$)
ViT-B-32-OpenAI 0.821 0.737 0.781
ViT-B-16-LAION 0.813 0.773 0.737
ViT-L-14-OpenAI 0.827 0.751 0.758
@@
Table 2: Results between human judgment and LLM judgment on CCS labelling. SC denotes Spearman’s correlation.
@@ 3 Quantifying interpretability in CLIP models
3.1 Preliminaries
In this section, we describe our methodology, starting with the TextSpan  (Gandelsman et al., ) algorithm and its extension across all attention heads in multiple CLIP models using in-context learning. TextSpan associates each attention head with relevant text descriptions by analyzing the variance in projections between head outputs and candidate text representations. Through iterative projections, it identifies distinct components aligned with different semantic aspects. While effective at linking heads to descriptive text spans, TextSpan does not assign explicit concept labels. In the next section, we detail our method for labeling the concepts learned by individual CLIP heads.
@@ 3.2.3 Evaluating Concept Consistency
To assess the consistency of a head with respect to its assigned concept label, we employ three state-of-the-art foundational models, GPT-4o, Gemini 1.5 pro and Claude Sonnet as external evaluators. For each TextSpan $T_{i}$ associated with head $h$, GPT-4o determines whether it aligns with the assigned concept $C_{h}$. The Concept Consistency Score (CCS) for head $h$ is then computed as:
@@
where $\mathbb{1}[\cdot]$ is an indicator function that returns 1 if $T_{i}$ to be consistent with $C_{h}$, and 0 otherwise. To ensure a high standard of reliability, we define consistency strictly—only if all three LLM judges independently rate $T_{i}$ as consistent with $C_{h}$. This requirement for unanimous agreement minimizes the influence of individual model biases or variability in judgment  (Liu & Zhang, 2025), thereby enhancing the robustness and trustworthiness of the overall concept consistency score.
@@
Next, we categorize each attention head based on its Concept Consistency Score (CCS) into three levels: high, moderate, and low. A head is considered to have a high CCS if all of its associated text descriptions align with the labeled concept, indicating that the head is highly specialized and likely encodes features relevant to that concept. Moderate CCS heads exhibit partial alignment, with three out of five text descriptions matching the concept label, suggesting that they capture the concept to some extent but not exclusively. In contrast, low CCS heads have zero or only one matching description, implying minimal relevance and indicating that these heads are largely unrelated to the given concept. This categorization provides insight into the degree of concept selectivity exhibited by individual attention heads. Table  1 shows examples of different types of CCS heads.
@@ 3.3 Evaluating LLM Judgment Alignment with Human Annotations
In the previous section, we introduced the Concept Consistency Score (CCS), computed using three LLM judges as an external evaluator. This raises an important question: Are LLM evaluations reliable and aligned with human assessments? To investigate this, we conducted a human evaluation study comparing LLM-generated judgments with human annotations. We selected 100 TextSpan descriptions from three different models, along with their assigned concept labels, and asked one of the authors to manually assess the semantic alignment between each span and its corresponding label.
Table 2 reports the agreement metrics between human and LLM evaluations, including Cohen’s Kappa, Spearman’s $\rho$, and Kendall’s $\uptau$. The Kappa values exceed 0.8, indicating extremely substantial agreement, while the correlation scores consistently surpass 0.7, confirming strong alignment. These results validate the use of LLMs as reliable evaluators in concept consistency analysis. The high agreement with human judgments suggests that LLMs can effectively assess semantic coherence, offering a scalable alternative to manual annotation. In the next section, we introduce the tasks and datasets used in our experiments.
@@
3.4 Experimental Setting
3.4.1 Tasks
Image classification: CIFAR-10  (Krizhevsky et al., 2009), CIFAR-100  (Krizhevsky et al., 2009), Food-101  (Bossard et al., 2014), Country-211  (Radford et al., 2021) and Oxford-pets  (Parkhi et al., 2012).
Out-of-domain classification: Imagenet-A  (Hendrycks et al., 2021b) and Imagenet-R  (Hendrycks et al., 2021a).
Video retrieval: MSRVTT  (Xu et al., 2016), MSVD  (Chen & Dolan, 2011), DiDeMo  (Anne Hendricks et al., 2017).
Bias: FairFace  (Karkkainen & Joo, 2021), SocialCounterFactuals  (Howard et al., 2024).
3.4.2 Models
For experiments we use the following six foundational image-text models: ViT-B-32, ViT-B-16 and ViT-L-14 pretrained from OpenAI-400M  (Radford et al., 2021) and LAION2B  (Schuhmann et al., 2022). Next, we discuss in detail the results from the experiments.
@@ 3.4.2 Models
Figure 2: Zero-shot performance comparison for CIFAR-10, CIFAR-100, and Food-101 datasets under different pruning strategies. For random pruning, results are averaged across five runs.
@@
Model Country-211 Oxford-pets ImageNet-A ImageNet-R
Original High CCS Low CCS Original High CCS Low CCS Original High CCS Low CCS Original High CCS Low CCS
ViT-B-32-OpenAI 17.16 11.46 16.3 50.07 46.66 48.96 31.49 20.24 28.72 69.09 54.47 64.45
ViT-B-32-datacomp 4.43 4.37 4.37 26.48 25.98 25.33 4.96 4.59 4.65 34.06 31.6 32.47
ViT-B-16-OpenAI 22.81 10.72 21.79 52.72 49.12 51.89 49.85 25.49 47.27 77.37 55.52 74.84
ViT-B-16-LAION 20.45 7.49 16.87 65.79 48.48 49.81 37.97 25.27 27.44 80.56 66.32 71.73
ViT-L-14-OpenAI 31.91 23.21 30.63 61.79 62.04 62.08 70.4 68.15 69.2 87.87 86.56 86.97
ViT-L-14-LAION 26.41 16.38 25.66 54.1 56.12 57.16 53.8 42.44 52.93 87.12 82.22 86.94
Table 4: Accuracy comparison of various CLIP models on Country-211, Oxford-pets, ImageNet-A and ImageNet-R datasets. The values represent original accuracy, performance after pruning high-CCS heads, and performance after pruning low-CCS heads.
4 Results and Discussion
4.1 Interpretable CLIP Models: The Role of CCS.
In this section we examine the role of the Concept Consistency Score (CCS) in revealing CLIP’s decision-making process, focusing on the question: How does CCS provide deeper insights into the functional role of individual attention heads in influencing downstream tasks? To explore this, we perform a soft-pruning analysis by zeroing out attention weights of heads with extreme CCS values—specifically, high CCS (CCS = 5) and low CCS (CCS $\leq$ 1). This approach disables selected heads without modifying the model architecture. As shown in Table  3, pruning high-CCS heads consistently causes significant drops in zero-shot classification performance across CIFAR-10, CIFAR-100 and FOOD-101 while pruning low-CCS heads has a minimal effect. This performance gap demonstrates that CCS effectively identifies heads encoding critical, concept-aligned information, making it a reliable tool for interpreting CLIP’s internal decision-making mechanisms.
We further observe notable variations in pruning sensitivity across model architectures. ViT-B-16 models suffer the most from high-CCS head pruning, implying a reliance on a smaller number of specialized heads. In contrast, ViT-L-14 models show greater resilience, suggesting more distributed representations. Among smaller models, OpenAI-trained models experience larger performance drops than OpenCLIP models when high-CCS heads are pruned. However, in larger models like ViT-L-14, OpenCLIP variants show a slightly higher degradation. These patterns reveal that CCS not only identifies functionally important heads but also captures model-specific and training-specific differences in how conceptual knowledge is organized and utilized within CLIP architectures.
4.2 High CCS vs random heads pruning
In the previous section, we showed that attention heads with high Concept Consistency Scores (CCS) are crucial to CLIP’s performance. To validate whether these heads are truly more important than others, we perform a controlled comparison against random pruning. Specifically, we randomly prune the same number of attention heads—excluding high-CCS heads—and repeat this across five seeds, averaging the results. As illustrated in Figure  2, pruning high-CCS heads consistently causes a significantly larger drop in zero-shot accuracy compared to random pruning across datasets and model variants. In contrast, random pruning results in only minor performance degradation, highlighting the functional importance of high-CCS heads. Interestingly, we also find that larger CLIP models show a smaller performance gap between high-CCS and random pruning, suggesting that larger architectures may be more robust due to greater redundancy or more distributed representations. These findings support CCS as a reliable and interpretable metric for identifying concept-relevant heads and offer deeper insights into how CLIP organizes conceptual information.